OpenStudy (anonymous):

The standard baseball measures 2 11/16" in diameter. What is its volume (to the nearest hundredth)? _____cu. in.

OpenStudy (anonymous):

@JakeV8

OpenStudy (anonymous):

what's the volume equation for a sphere?

OpenStudy (anonymous):

4/3pir^3

OpenStudy (anonymous):

that's better :) So find the radius...

OpenStudy (anonymous):

we gonna divide

OpenStudy (anonymous):

2 11/16

OpenStudy (anonymous):

that's the diameter, so you need to divide 2 11/16 in half to get the radius.. .careful with the fractions :)

OpenStudy (anonymous):

0.6875=11/16

OpenStudy (anonymous):

43

OpenStudy (anonymous):

diameter = 2 11/16 = 2.6875 radius = diameter / 2 = 1.34375 Volume = (4/3) pi (radius^3) = (4/3) pi (1.34375^3) = (4/3) pi (2.4264) = 5.8416 pi = 18.3519 =18.35 cm^3 <<-- rounded to nearest hundredths That's what I got... not sure where you got 43? But I could have made an error... might want to check it.